OpenStudy (anonymous):

what is the following expression equal _____ √ 62-2i^2 _____ A.)√60i B.)8i C.) 8 ____ ____ D.) √62 -√ 2i

OpenStudy (anonymous):

Please don't post multiple questions all at once. You won't be able to pay attention if two people are trying to help you, and it crowds out questions other people may have.

OpenStudy (anonymous):

That being said, is the original expression a)\((\sqrt{62-2i})^2\) b)\(\sqrt{62-2i^2}\) c)None of the above.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

\(i^2 = -1 \implies -2i^2 = -2(-1) = 2\)

OpenStudy (anonymous):

So what would you have under the radical?

OpenStudy (anonymous):

i think the answer is c

OpenStudy (anonymous):

Yes, because you have \(\sqrt{64}\)